When 40 kV is applied across an X-ray tube, X-ray is obtained with a maximum frequency of 9.7 × 10^18 Hz. Calculate the value of Planck constant from these data.

Grade:11

1 Answers

Kevin Nash
9 years ago
λ = hc/eV ; V = 40 kV, f = 9.7 * 10^18 Hz Or, h/c = h/eV ; or, i/f = h/eV ; or h = eV/f V – s = eV/f V –s = 40* 10^3/9.7* 10^18 = 4.12 * 10^15 eV-s.